OpenStudy (anonymous):

which of the following statements is (are) true for all postitive intergers 1) 2+2^2+2^3+....+2n=2n+1-2 2) 1+3+5+...+n=n^3

OpenStudy (anonymous):

1) i think there needs to be exponents to understand it correctly? Is this correct? \[2+2^2+2^3+\cdots+2^n=2^{n+1}-2\] 2)as written the formula doesnt make sense, should it be this? \[1+3+5+\cdots+(2n-1)=n^3\]The reason for the 2n-1 is because you only want odd numbers.

OpenStudy (anonymous):

ohh i see cool thanks man :)